# Modelling conductive thermal transport in three-dimensional fibrous media with fiber-to-fiber contacts

### v1

@@ Abstract
Understanding heat transfers in fibrous materials, particularly conduction, is a major challenge due to their heterogeneous and multiscale nature, and the unknown contribution of fiber-to-fiber contacts. In most previous modelling studies, the existence of thermal contact resistance is not considered, and the computational complexity often limits the size of simulated samples, which often leads to imprecise of inaccurate predictions. The same problem arises when considering electrical conduction through fibrous materials. In this work, we describe a computationally efficient simulation approach based on multi-nodal representation to analyze the steady-state heat conduction through the solid structure in numerically generated 3D fibrous networks, including contact resistance. We show that the solid conductivity in these networks is governed by a master curve that depends on a single parameter: a characteristic ratio representing the interplay between fiber intrinsic conductivity and contact resistance as well as the influence of other geometric parameters, which numerically validates previous theoretical studies. However, we observe a deviation to this established theory for poorly-connected networks. We derive an expression for a correction factor, considering the influence of correlations between fiber temperatures, and we find then good agreement with our simulation data. Our results demonstrate that the solid conductivity can be fully predicted based on geometric quantities, regardless of the extent of network connectivity, thus generalizing previous studies on this topic. This work, contributing to improve our understanding of conductive heat transport in fibrous media, may prove useful in the development of accurate predictive models and optimization strategies for fibrous insulation materials.
@@ I Introduction
Reducing energy costs in the building sector drives important efforts to improve the performance of insulation materials. These materials are generally porous media, in the form of foams or fibrous assemblies. In particular, glasswool, which is composed of glass fibers dispersed in air and linked together by a binder, is one of the most common insulation currently on the market. The thermal performance of an insulation material can be quantified by its effective thermal conductivity $k_{eff}$. As this quantity refers to a macroscopic scale, it accounts for multiple heat transfer mechanisms: convection, conduction through gas, conduction through the solid phase, and radiation. Improving the efficiency of fibrous insulations requires a quantitative knowledge of these four contributions to heat transport, and more precisely of all the parameters influencing it at the microscale. In media with a pore size below 1 cm, which is the case of glasswool, convection can be neglected [1], leaving only the contributions of conduction and radiation, which can be respectively represented by two independent terms $k_{cond}$ and $k_{rad}$. Under the hypothesis of low-density, below about 150 kg.m-3, which is applicable to glasswool, the conductive component can be split as a sum of an air conductivity component $k_{cond,gas}$ and a solid conductivity component $k_{cond,solid}$. This results in the following additive approximation [2]: $k_{eff}=k_{rad}+k_{cond,gas}+k_{cond,solid}$, which means that the three identified contributions to heat transfer are assumed to be uncoupled and can be studied separately. Although many studies have been conducted to quantify the radiative contribution in fibrous materials [3, 4, 5, 6], only a few have considered the conduction term and its sensitivity to parameter changes. As the contribution of the gas conductivity can be considered as constant with medium’s density and microstructure, it is interesting to further investigate the impact of conduction through the solid phase in fibrous insulation materials, with the aim of understanding and, subsequently, minimizing this contribution.
As the separate contributions to the effective conductivity cannot be easily measured experimentally, numerical modelling is the preferred tool for their quantitative analysis. The thermal performance of fibrous materials can be investigated numerically using either real structures reconstructed from images, that can be acquired by X-ray tomography [7, 8, 9], or virtual samples generated by algorithms [10, 11, 12]. The last option enables artificial media to be fitted to real materials using geometric parameters extracted from tomography images, without considering the whole reconstructed 3D structure for the heat transfer calculations, thus improving computational efficiency.
@@
The case of fibrous insulation materials such as mineral wools has been previously studied by Chapelle et al. [22], Arambakam et al. [26] and Kallel and Joulain [27]. In order to be as representative as possible in terms of network geometry, the authors have chosen a given fiber orientation distribution, first introduced by Schladitz et al. [28] for a stacked fiber system. Among its characteristic is an invariance along the $x$ and $y$ axes, representing the morphology of fibrous insulation materials in relation to their manufacturing process. In these studies, a modified random walk generation method has been used for each fiber, allowing the successive positions of the spheres to be controlled by two parameters so as to create tortuosity.
@@
The volume of the 3D fibrous medium that has to be numerically generated should be set to exceed the representative volume element (RVE), defined as the smallest simulation volume for which the computed properties - in this case the solid conductivity, will be representative of the whole material. This solid conductivity can be numerically computed by different methods. Finite element analysis, used in Refs. [27] and [8], allows a complete representation of the studied structures, resulting in precise yet computationally demanding simulations of the heat transfers, in particular due to the meshing step. An alternative approach to reduce computation time and thus increase the simulation volumes that can be reached is the representation of a fibrous network by a circuit of nodes and branches. This procedure, first used by Kirkpatrick to simulate electrical conductivity [29], is only applicable to the separate calculation of the contribution of conduction through the solid phase, as opposed to volume averaging methods [30]. Although most studies exploiting this method have applied it to quantifying the solid electrical conductivity of metallic nanotube networks [31, 32, 33], very few have extended it to the resolution of thermal transfer problems so far [34].
@@
Independently of the chosen methods for fiber generation and heat transfer calculation in fibrous media, the dependence of thermal or electrical solid conductivity on network parameters has been numerically investigated in various studies [27, 35, 36], only some of which have included the effect of fiber-to-fiber contact resistance [37, 38, 39, 40]. It has been found in particular that the Biot number, which compares the conductive thermal resistance in the fiber with the external surface resistance due to thermal losses through contacts [41], has a noticeable impact on the thermal conductivity of fibrous materials [42]. The reported studies have suggested, in some cases, simplified theoretical models [38, 43, 44, 45], yet constrained to specific limiting scenarios. Recently, a statistical analysis combined with Fourier’s law has been used to derive a more general theoretical model that predicts the thermal conductivity of dense networks of straight nanofibers, in 2D by Volkov and Zhigilei [46, 47], and in 2D and 3D by Zhao et al. [48]. This theoretical framework generalizes the results of the above-mentioned work by introducing the dependence of solid conductivity on a geometry-dependent resistance ratio. In these studies, the outputs of the model have been exclusively compared with results of numerical simulations in the case of 2D carbon nanotube networks; it is only very recently that Volkov and Zhigilei [49] have presented a first comparison of 3D simulation results, providing at the same time a more accurate theoretical description of semi-dilute systems. However, to the best of our knowledge, there have been no reported studies conducting numerical simulations of 3D networks with variations of all fiber geometric parameters and of contact resistance, making the virtual media representative of a broad spectrum of fibrous materials instead of focusing on an application to metallic nanowires. In particular, anisotropic and poorly-connected 3D systems, which represent the typical case of insulation materials, have not been extensively investigated so far, especially theoretically.
@@
In this work, we focus on quantifying the separate influence of the contribution of conduction through the solid phase to the total heat transfer in fibrous insulation materials, which also includes radiation and conduction through the gas phase. We use a multi-nodal representation to innovatively study the solid conductivity in virtually generated 3D media with straight fibers, with the specific intention to apply the results to the case of real glasswool samples. A methodological advantage of this method is the possibility of directly integrating a fiber-to-fiber contact resistance as an adjustable parameter, while maintaining reasonable computational efficiency. We demonstrate in particular that the impact of fiber parameters - orientation, length and diameter, as well as that of density, on the conductive properties strongly depends on this contact resistance. More generally, we present a novel approach to quantify the evolution of the solid conductivity in 3D fibrous networks under the form of a master curve, valid regardless of the individual value of these four fiber parameters. We obtain good agreement between our simulation results and the theoretical model introduced by Zhao et al. [48] and Volkov and Zhigilei [46, 47] for high fiber aspect ratios or high fiber volume fractions, and we show that an adaptation of the model is needed in other cases, in accordance with the latest work from Volkov and Zhigilei [49]. A first novelty of our work lies in the validation of this modified theoretical model with simulation data, showing its accuracy for the prediction of the solid conductivity for a wide range of 3D networks including anisotropic and poorly-connected ones, which goes beyond the commonly studied case of metallic nanowires assemblies. Another innovative discovery is the simple formulation of this model depending on geometrical parameters only, revealing in particular the significance of the average number of contacts per fiber as a key parameter governing the extent of the needed correction compared with the initial high-density case. Finally, we have shown a comparison of our simulation and theoretical results, focusing on the specific case of glasswool. This approach gives first insights into the potential impact of contacts on thermal performance at product scale. We also emphasize that our numerical modeling is well suited to predict the electrical conductivity of fibrous media.
@@
This article is organized as follows. In section II.1, we explain the fibrous medium generation method used in this work. Then, in section II.2, we describe the nodal method used to represent the generated networks by linear circuits and to compute the steady-state heat conduction through the solid phase. Next, numerical results are reported and discussed in section III; the effect of fiber parameters and contact resistance on the evolution of solid conductivity in 3D fibrous networks is analyzed, theoretical considerations are introduced based on simulation results. The final section concludes the paper and outlines interesting research directions for future work.
@@ II.1 Virtual fibrous medium generation
The generated medium is a three-dimensional assembly of straight lines, representing fibers, with an orientation distribution and a prescribed length $l$. This choice is motivated by the low tortuosity of fibers in glasswool as shown by Optical Coherence Tomography acquisitions, an example of which is provided in Fig. S1 (Supplementary Material) [50]. For each line, one extremity is randomly placed in a cubic computational domain of side $L$. Its orientation is defined by a pair of angles ($\theta_{i}$,$\varphi_{i}$) in spherical coordinates, as shown in Fig. 1
Figure 1: Examples of generated 3D networks, in which the orientation of one fiber $i$ is described by a polar angle $\theta_{i}$ and an azimuthal angle $\varphi_{i}$ in the three-dimensional space (a). In these simulations, $L=3$ mm, the final volume fraction reached is $V_{f}=0.0033$, fiber diameter is $d=10~{}\mu$m, fiber length is $l=3$ mm and the average polar angle is (b) $\bar{\theta}=31.8^{\circ}$ ($\beta=0.1$), (c) $\bar{\theta}=57.3^{\circ}$ ($\beta=1$), and (d) $\bar{\theta}=79.9^{\circ}$ ($\beta=8$). The yellow (resp. red) circles represent the contact points between a fiber and the upper (resp. lower) edge of the simulation volume, and the blue circles represent the detected contact points between fibers.
(a), sampled according to a directional distribution, the $\beta$-orientation distribution, previously used in different modelling studies [28, 21, 26, 22, 27].
The azimuthal $\varphi_{i}$ angle is chosen from a uniform distribution between 00 and $2\pi$: the density of the distribution is independent of $\varphi$, representing the isotropy of the material in the $xy$-plane. For the choice of the polar angle $\theta_{i}$, the following marginal probability density function is applied: $g_{\beta}(\theta)=\frac{\beta\sin\theta}{2[1+(\beta^{2}-1)\cos^{2}\beta]^{3/2}},\theta\in[0,\pi],\beta\in R^{+}\{0\}$. The adjustable parameter $\beta$, called the anisotropy parameter, controls the shape and the mean value of the distribution. The case $\beta=1$ describes an isotropic system; for $\beta\rightarrow 0$, the distribution concentrates around the $z$-axis and for $\beta\rightarrow\infty$, the orientations are isotropically distributed in the $xy$-plane. For the sake of convenience, the average polar angle of all fibers $\bar{\theta}$ will be referred to as the network parameter for fiber orientation, rather than $\beta$.
@@
The number of lines initially generated in the system is chosen according to a preimposed volume fraction, called the initial volume fraction, and considering them as cylinders with a fixed value $d$ for their diameter. The contacts are then identified using a distance threshold: two lines are considered connected if the smallest distance between them is inferior to their diameter $d$. In this approximation, the connected fibers are modelled by soft-core, i.e., interpenetrating objects. The soft-core approach does not consider correlations in the spatial distribution of lines, as they may exist in a physical system of connected fibers. However, it has been reported in previous studies [24, 52] that the error introduced by this approximation is small, less than 1% for an aspect ratio $a=l/d\geq 100$, which is the average one in glasswool [53]. It should be noted that lines having less than two contact points cannot be included in the nodal analysis described in the following, as the resulting thermal network has to be fully connected to be solved by matrix approach. As they do not contribute to heat transfer, after contact detection, these lines are removed from the generated medium, as well as the associated contact points through an iterative procedure. In this study, the volume fraction of the medium $V_{f}$ is defined as the final one, obtained after fiber removal and therefore inferior to the initial one. Figs. 1(b), (c) and (d) show generated networks using our numerical approach with identification of contact points, for three different angle distributions controlled by the parameter $\beta$.
@@ II.2 Nodal analysis for the resolution of a steady-state conductive heat transport problem
The steady-state conduction through the solid phase in the generated network is computed numerically using the configuration of the guarded hot plate experiment. The upper and lower boundaries of the simulation domain represent two plates with different temperatures, which means that each fiber extremity located at $z=0$ (resp. $z=L$) will be a temperature source at $T=T_{b}$ (resp. $T=T_{b}+\Delta T$), while the fibers will be considered as conductive with an intrinsic thermal conductivity $k_{fib}=1.3$ Wm-1K-1. This value corresponds to the thermal conductivity of silica glass, the fiber material in glasswool mats. The four other boundaries, on the side of the simulation domain, are assumed to be adiabatic. In the isotropic case, the solid thermal conductivity is the same in each direction. However, if the anisotropy parameter of the distribution orientation is different from unity, the transverse component (i.e., the thermal conductivity in the $z$ direction) differs from the planar component (i.e., that in the $x$ or $y$ directions). This is particularly the case for fibrous insulation materials, due to their layered fiber arrangement [54]. Note that here, in the guarded hot plate experiment configuration, we exclusively simulate the transverse solid thermal conductivity, referred to in this work as the solid conductivity, meaning the heat flux direction will always be the cross-plane direction for the fibrous medium.
@@
The generated network has to be sufficiently dense or large to reach the RVE and to allow for a low-incertitude prediction of the solid conductivity. By analogy with an electrical circuit, it can be computed using graph theory, with the Modified Nodal Analysis method [32]. In the steady state, this thermal network is a linear circuit arranged in branches, representing heat flow rates crossing thermal resistances, connected to each other by nodes representing temperatures or temperature sources [34, 55]. This nodal method assumes uniform sections in temperature, i.e., the conductive heat transfer in each fiber will be considered as unidimensional, occurring only in the direction of the fiber. This choice is justified by the average aspect ratio of fibers $a\approx 100$ in glasswool [53]. In particular, the network is modelled using multi-nodal representation, first introduced by Da Rocha et al. [31] for electrical circuits, considering both the internal thermal resistance of fibers and the thermal contact resistance. This modelling method is represented in Fig. 2
@@
Figure 2: (a) Simplified network of three connected fibers. Contacts are identified with blue circles. (b) Equivalent thermal circuit of resistances in the steady-state configuration. The dots represent nodes, in particular, yellow and red dots corresponds to temperature sources. A given branch $f$ can be associated with either a contact resistance $R_{k}$, or an internal resistance $R_{w}l_{f}$ proportional to its length where $R_{w}=\frac{4}{k_{fib}\pi d^{2}}$.
@@
. In this kind of model, each node represents a contact point: either between a fiber and a temperature source on the upper or lower edge of the simulation cube, or between two fibers. In the latter case, the junction is defined by two distinct nodes linked by an additional branch, that represents the fiber-to-fiber contact resistance. This kind of model differs from the junction resistance-dominated assumption, in which one node represents one fiber, linked by contact resistances, thus ignoring the contribution of the internal resistances. Also, as it can be seen in Fig. 2, when fiber extremities are not touching any other fiber, they are assumed to be adiabatic and therefore are not included in the thermal network.
@@
As the considered heat transfer problem is a steady-state configuration, the associated thermal network can be described by algebraic graph theory. The resolution of a single matrix equation allows to get the vector of temperatures and that of heat fluxes, making possible the representation of the temperature field in the network as shown in Fig. 3
@@
Figure 3: Temperature field distributions in generated networks with $L=3$ mm, $V_{f}=0.0198$, $l=2$ mm, $d=10~{}\mu$m and (a) $\bar{\theta}=31.8^{\circ}$ ($\beta=0.1$), (b) $\bar{\theta}=57.3^{\circ}$ ($\beta=1$), and (c) $\bar{\theta}=79.9^{\circ}$ ($\beta=8$), containing between 3720 and 3990 fibers. The upper and lower edges of the simulation domain are areas where the temperature is imposed, generating a heat flux $Q$ in the vertical direction. The angle $\theta_{i}$ of each fiber $i$ is therefore its angle with respect to the direction of heat flux. The coordinate system used is the one represented in Fig. 1(a).
@@
for different angle distributions (see Appendix A for details on the resolution method). Then, the total heat flux $Q$ is computed by adding the contributions from the branches in contact with the upper or the lower edge of the simulation domain, and the solid thermal conductivity is calculated by: $k_{cond,solid}=QL/(L^{2}\Delta T)$. As the thermal properties of individual fibers, the fiber conductivity $k_{fib}$ and the fiber-to-fiber contact resistance $R_{k}$, are assumed to be independent of the temperature, the solid conductivity does not depend on $T_{b}$ and $\Delta T$.

### v2

# Modeling conductive thermal transport in three-dimensional fibrous media with fiber-to-fiber contacts

@@ Abstract
Understanding heat transfers in fibrous materials, particularly conduction, is a major challenge due to their heterogeneous and multiscale nature, and the unknown contribution of fiber-to-fiber contacts. In most previous modeling studies, the existence of thermal contact resistance is not considered, and the computational complexity limits the size of simulated samples, which often leads to imprecise or inaccurate predictions. The same problem arises when considering electrical conduction through fibrous materials. In this work, we describe a computationally efficient simulation approach based on multinodal representation to analyze the steady-state heat conduction through the solid structure in numerically generated three-dimensional fibrous networks, including contact resistance. We show that the solid conductivity in these networks is governed by a master curve that depends on a single parameter: a characteristic ratio representing the interplay between the intrinsic fiber conductivity and contact resistance as well as the influence of other geometric parameters, which numerically validates previous theoretical studies. However, we observe a deviation to this established theory for poorly connected networks. We derive an expression for a correction factor, considering the influence of correlations between fiber temperatures, and we then find good agreement with our simulation data. Our results demonstrate that the solid conductivity can be fully predicted based on geometric quantities, regardless of the extent of network connectivity, thus generalizing previous studies on this topic. This work, contributing to improve our understanding of conductive heat transport in fibrous media, may prove useful in the development of accurate predictive models and optimization strategies for fibrous insulation materials.
@@ I Introduction
Reducing energy costs in the building sector drives important efforts to improve the performance of insulation materials. These materials are generally porous media, in the form of foams or fibrous assemblies. In particular, glasswool, which is composed of glass fibers dispersed in air and linked together by a binder, is one of the most common insulations currently on the market. The thermal performance of an insulation material can be quantified by its effective thermal conductivity $k_{eff}$. As this quantity refers to a macroscopic scale, it accounts for multiple heat transfer mechanisms: convection, conduction through gas, conduction through the solid phase, and radiation. Improving the efficiency of fibrous insulations requires a quantitative knowledge of these four contributions to heat transport, and more precisely of all the parameters influencing it at the microscale. In media with a pore size below 1 cm, which is the case for glasswool, convection can be neglected [1], leaving only the contributions of conduction and radiation, which can be respectively represented by two independent terms $k_{cond}$ and $k_{rad}$. Under the hypothesis of low density, below about 150 kg m-3, which is applicable to glasswool, the conductive component can be split as a sum of an air conductivity component $k_{cond,gas}$ and a solid conductivity component $k_{cond,solid}$. This results in the additive approximation [2] $k_{eff}=k_{rad}+k_{cond,gas}+k_{cond,solid}$, which means that the three identified contributions to heat transfer are assumed to be uncoupled and can be studied separately. Although many studies have been conducted to quantify the radiative contribution in fibrous materials [3, 4, 5, 6], only a few have considered the conduction term and its sensitivity to parameter changes. As the contribution of gas conductivity can be considered as constant with the medium’s density and microstructure, it is interesting to further investigate the impact of conduction through the solid phase in fibrous insulation materials, with the aim of understanding and, subsequently, minimizing this contribution.
As the separate contributions to the effective conductivity cannot be easily measured experimentally, numerical modeling is the preferred tool for their quantitative analysis. The thermal performance of fibrous materials can be investigated numerically using either real structures reconstructed from images, that can be acquired by X-ray tomography [7, 8, 9], or virtual samples generated by algorithms [10, 11, 12]. The last option enables artificial media to be fitted to real materials using geometric parameters extracted from tomography images, without considering the whole reconstructed three-dimensional (3D) structure for the heat transfer calculations, thus improving computational efficiency.
@@
The case of fibrous insulation materials such as mineral wools has been previously studied by Chapelle et al. [22], Arambakam et al. [26], and Kallel and Joulain [27]. In order to be as representative as possible in terms of network geometry, the authors chose a given fiber orientation distribution, first introduced by Schladitz et al. [28] for a stacked fiber system. Among its characteristic is an invariance along the $x$ and $y$ axes, representing the morphology of fibrous insulation materials in relation to their manufacturing process. In these studies, a modified random walk generation method was used for each fiber, allowing the successive positions of the spheres to be controlled by two parameters so as to create tortuosity.
@@
The volume of the 3D fibrous medium that has to be numerically generated should be set to exceed the representative volume element (RVE), defined as the smallest simulation volume for which the computed properties - in this case the solid conductivity - will be representative of the whole material. This solid conductivity can be numerically computed by different methods. Finite-element analysis, used in Refs. [27] and [8], allows a complete representation of the studied structures, resulting in precise yet computationally demanding simulations of the heat transfers, in particular due to the meshing step. An alternative approach to reduce computation time and thus increase the simulation volumes that can be reached is the representation of a fibrous network by a circuit of nodes and branches. This procedure, first used by Kirkpatrick [29] to simulate electrical conductivity, is only applicable to the separate calculation of the contribution of conduction through the solid phase, as opposed to volume averaging methods [30]. Although most studies exploiting this method have applied it to quantifying the solid electrical conductivity of metallic nanotube networks [31, 32, 33], very few have extended it to the resolution of thermal transfer problems so far [34].
@@
Independently of the chosen methods for fiber generation and heat transfer calculation in fibrous media, the dependence of thermal or electrical solid conductivity on network parameters has been numerically investigated in various studies [27, 35, 36], only some of which have included the effect of fiber-to-fiber contact resistance [37, 38, 39, 40]. It has been found in particular that the Biot number, which compares the conductive thermal resistance in the fiber with the external surface resistance due to thermal losses through contacts [41], has a noticeable impact on the thermal conductivity of fibrous materials [42]. The reported studies have suggested, in some cases, simplified theoretical models [38, 43, 44, 45], yet constrained to specific limiting scenarios. Recently, a statistical analysis combined with Fourier’s law has been used to derive a more general theoretical model that predicts the thermal conductivity of dense networks of straight nanofibers, in two dimensions by Volkov and Zhigilei [46, 47], and in two and three dimensions by Zhao et al. [48]. This theoretical framework generalizes the results of the abovementioned work by introducing the dependence of solid conductivity on a geometry-dependent resistance ratio. In these studies, the outputs of the model have been exclusively compared with results of numerical simulations in the case of 2D carbon nanotube networks; it is only recently that Volkov and Zhigilei [49] presented a first comparison of 3D simulation results, providing at the same time a more accurate theoretical description of semidilute systems. However, to the best of our knowledge, there have been no reported studies conducting numerical simulations of 3D networks with variations of all fiber geometric parameters and of contact resistance, making the virtual media representative of a broad spectrum of fibrous materials instead of focusing on an application to metallic nanowires. In particular, anisotropic and poorly connected 3D systems, which represent the typical case of insulation materials, have not been extensively investigated so far, especially theoretically.
@@
In this work, we focus on quantifying the separate influence of the contribution of conduction through the solid phase to the total heat transfer in fibrous insulation materials, which also includes radiation and conduction through the gas phase. We use a multinodal representation to innovatively study the solid conductivity in virtually generated 3D media with straight fibers, with the specific intention to apply the results to the case of real glasswool samples. A methodological advantage of this method is the possibility of directly integrating a fiber-to-fiber contact resistance as an adjustable parameter, while maintaining reasonable computational efficiency. We demonstrate in particular that the impact of fiber parameters - orientation, length and diameter - as well as that of density, on the conductive properties strongly depends on this contact resistance. More generally, we present an approach to quantify the evolution of the solid conductivity in 3D fibrous networks under the form of a master curve, valid regardless of the individual value of these four fiber parameters. We obtain good agreement between our simulation results and the theoretical model introduced by Zhao et al. [48] and Volkov and Zhigilei [46, 47] for high fiber aspect ratios or high fiber volume fractions, and we show that an adaptation of the model is needed in other cases, in accordance with the latest work from Volkov and Zhigilei [49]. A first contribution of our work lies in the validation of this modified theoretical model with simulation data, showing its accuracy for the prediction of the solid conductivity for a wide range of 3D networks including anisotropic and poorly connected ones, which goes beyond the commonly studied case of metallic nanowire assemblies. Another innovative discovery is the simple formulation of this model depending on geometrical parameters only, revealing in particular the significance of the average number of contacts per fiber as a key parameter governing the extent of the needed correction compared with the initial high-density case. Finally, we show a comparison between our simulation and theoretical results, focusing on the specific case of glasswool. This approach gives first insights into the potential impact of contacts on thermal performance at the product scale. We also emphasize that our numerical modeling is well suited to predict the electrical conductivity of fibrous media.
@@
This article is organized as follows. In section II.1, we explain the fibrous medium generation method used in this work. Then, in section II.2, we describe the nodal method used to represent the generated networks by linear circuits and to compute the steady-state heat conduction through the solid phase. Next, numerical results are reported and discussed in section III; the effect of fiber parameters and contact resistance on the evolution of solid conductivity in 3D fibrous networks is analyzed, and theoretical considerations are introduced based on simulation results. The final section concludes the paper and outlines interesting research directions for future work.
@@ II.1 Virtual fibrous medium generation
The generated medium is a three-dimensional assembly of straight lines, representing fibers, with an orientation distribution and a prescribed length $l$. This choice is motivated by the low tortuosity of fibers in glasswool as shown by optical coherence tomography acquisitions, an example of which is provided within the Supplemental Material [50]. For each line, one extremity is randomly placed in a cubic computational domain of side $L$. Its orientation is defined by a pair of angles ($\theta_{i}$,$\varphi_{i}$) in spherical coordinates, as shown in Fig. 1
Figure 1: Examples of generated 3D networks, in which the orientation of one fiber $i$ is described by a polar angle $\theta_{i}$ and an azimuthal angle $\varphi_{i}$ in the three-dimensional space (a). In these simulations, $L=3$ mm, the final volume fraction reached is $V_{f}=0.0033$, the fiber diameter is $d=10~{}\mu$m, the fiber length is $l=3$ mm and the average polar angle is (b) $\bar{\theta}=31.8^{\circ}$ ($\beta=0.1$), (c) $\bar{\theta}=57.3^{\circ}$ ($\beta=1$), and (d) $\bar{\theta}=79.9^{\circ}$ ($\beta=8$). The yellow (respectively red) circles represent the contact points between a fiber and the upper (respectively lower) edge of the simulation volume, and the blue circles represent the detected contact points between fibers.
(a), sampled according to a directional distribution, the $\beta$-orientation distribution, previously used in different modeling studies [28, 21, 26, 22, 27].
The azimuthal $\varphi_{i}$ angle is chosen from a uniform distribution between 00 and $2\pi$: the density of the distribution is independent of $\varphi$, representing the isotropy of the material in the $x$-$y$ plane. For the choice of the polar angle $\theta_{i}$, the following marginal probability density function is applied: $g_{\beta}(\theta)=\frac{\beta\sin\theta}{2[1+(\beta^{2}-1)\cos^{2}\beta]^{3/2}},\theta\in[0,\pi],\beta\in R^{+}\{0\}$. The adjustable parameter $\beta$, called the anisotropy parameter, controls the shape and the mean value of the distribution. The $\beta=1$ case describes an isotropic system; for $\beta\rightarrow 0$, the distribution concentrates around the $z$ axis and for $\beta\rightarrow\infty$, the orientations are isotropically distributed in the $x$-$y$ plane. For the sake of convenience, the average polar angle of all fibers $\bar{\theta}$ will be referred to as the network parameter for fiber orientation, rather than $\beta$.
@@
The number of lines initially generated in the system is chosen according to a preimposed volume fraction, called the initial volume fraction, and considering them as cylinders with a fixed value $d$ for their diameter. The contacts are then identified using a distance threshold: two lines are considered connected if the smallest distance between them is inferior to their diameter $d$. In this approximation, the connected fibers are modeled by soft-core, i.e., interpenetrating, objects. The soft-core approach does not consider correlations in the spatial distribution of lines, as they may exist in a physical system of connected fibers. However, it has been reported in previous studies [24, 52] that the error introduced by this approximation is small, less than 1% for an aspect ratio $a=l/d\geq 100$, which is the average one in glasswool [53]. It should be noted that lines having less than two contact points cannot be included in the nodal analysis described in the following, as the resulting thermal network has to be fully connected to be solved by the matrix approach. As they do not contribute to heat transfer, after contact detection, these lines are removed from the generated medium, as well as the associated contact points through an iterative procedure. In this study, the volume fraction of medium $V_{f}$ is defined as the final one, obtained after fiber removal and therefore inferior to the initial one. Figures 1(b), 1(c) and 1(d) show generated networks using our numerical approach with identification of contact points, for three different angle distributions controlled by parameter $\beta$.
@@ II.2 Nodal analysis for the resolution of a steady-state conductive heat transport problem
The steady-state conduction through the solid phase in the generated network is computed numerically using the configuration of the guarded hot plate experiment. The upper and lower boundaries of the simulation domain represent two plates with different temperatures, which means that each fiber extremity located at $z=0$ (respectively $z=L$) will be a temperature source at $T=T_{b}$ (respectively $T=T_{b}+\Delta T$), while the fibers will be considered as conductive with an intrinsic thermal conductivity $k_{fib}=1.3$ Wm-1K-1. This value corresponds to the thermal conductivity of silica glass, the fiber material in glasswool mats. The four other boundaries, on the side of the simulation domain, are assumed to be adiabatic. In the isotropic case, the solid thermal conductivity is the same in each direction. However, if the anisotropy parameter of the distribution orientation is different from unity, the transverse component (i.e., the thermal conductivity in the $z$ direction) differs from the planar component (i.e., that in the $x$ or $y$ direction). This is particularly the case for fibrous insulation materials, due to their layered fiber arrangement [54]. Note that here, in the guarded hot plate experiment configuration, we exclusively simulate the transverse solid thermal conductivity, referred to in this work as the solid conductivity, meaning that the heat flux direction will always be the cross-plane direction for the fibrous medium.
@@
The generated network has to be sufficiently dense or large to reach the RVE and to allow for a low-incertitude prediction of the solid conductivity. By analogy with an electrical circuit, it can be computed using graph theory, with the modified nodal analysis method [32]. In the steady state, this thermal network is a linear circuit arranged in branches, representing heat flow rates crossing thermal resistances, connected to each other by nodes representing temperatures or temperature sources [34, 55]. This nodal method assumes uniform sections in temperature, i.e., the conductive heat transfer in each fiber will be considered as unidimensional, occurring only in the direction of the fiber. This choice is justified by the average aspect ratio of fibers $a\approx 100$ in glasswool [53]. In particular, the network is modeled using a multinodal representation, first introduced by Da Rocha et al. [31] for electrical circuits, considering both the internal thermal resistance of fibers and the thermal contact resistance. This modeling method is represented in Fig. 2
@@
Figure 2: (a) Simplified network of three connected fibers. Contacts are identified with blue circles. (b) Equivalent thermal circuit of resistances in the steady-state configuration. The dots represent nodes; in particular, yellow and red dots correspond to temperature sources. A given branch $f$ can be associated with either a contact resistance $R_{k}$ or an internal resistance $R_{w}l_{f}$ proportional to its length, where $R_{w}=\frac{4}{k_{fib}\pi d^{2}}$.
@@
. In this kind of model, each node represents a contact point: either between a fiber and a temperature source on the upper or lower edge of the simulation cube, or between two fibers. In the latter case, the junction is defined by two distinct nodes linked by an additional branch that represents the fiber-to-fiber contact resistance. This kind of model differs from the junction resistance-dominated assumption, in which one node represents one fiber, linked by contact resistances, thus ignoring the contribution of the internal resistances. Also, as can be seen in Fig. 2, when fiber extremities are not touching any other fiber, they are assumed to be adiabatic and are therefore not included in the thermal network.
@@
As the considered heat transfer problem is a steady-state configuration, the associated thermal network can be described by algebraic graph theory. The resolution of a single matrix equation allows us to get the vector of temperatures and that of heat fluxes, making possible the representation of the temperature field in the network, as shown in Fig. 3
@@
Figure 3: Temperature field distributions in generated networks with $L=3$ mm, $V_{f}=0.0198$, $l=2$ mm, $d=10~{}\mu$m, and (a) $\bar{\theta}=31.8^{\circ}$ ($\beta=0.1$), (b) $\bar{\theta}=57.3^{\circ}$ ($\beta=1$), and (c) $\bar{\theta}=79.9^{\circ}$ ($\beta=8$), containing between 3720 and 3990 fibers. The upper and lower edges of the simulation domain are areas where the temperature is imposed, generating a heat flux $Q$ in the vertical direction. The angle $\theta_{i}$ of each fiber $i$ is therefore its angle with respect to the direction of the heat flux. The coordinate system used is that represented in Fig. 1(a).
@@
for different angle distributions (see Appendix A for details on the resolution method). Then, the total heat flux $Q$ is computed by adding the contributions from the branches in contact with the upper or lower edge of the simulation domain, and the solid thermal conductivity is calculated as $k_{cond,solid}=QL/(L^{2}\Delta T)$. As the thermal properties of individual fibers, the fiber conductivity $k_{fib}$ and the fiber-to-fiber contact resistance $R_{k}$, are assumed to be independent of the temperature, the solid conductivity does not depend on $T_{b}$ and $\Delta T$.